# First Workshop on LLM Persona Modeling

### v1

@@ 1 Introduction
Modern dating platforms primarily rely on profile similarity metrics, yet decades of relationship research show that compatibility emerges from interaction dynamics, not static attributes finkel2012online. We propose Love First, Know Later: instead of comparing static profiles to predict compatibility, we simulate the relationship itself first.
Recent advances show that LLMs can express consistent personality traits park2023generative; li2024big5chat; wang2025incharacter, follow human psychological patterns kosinski2023theory, and engage in social interactions yang2024oasis; zhou2024sotopia; liu2024roleagent. These models encode rich social priors from their training data, enabling them to simulate human behaviors and emotional responses. We leverage these capabilities to use LLMs as proxies for simulating romantic interactions. Building on the finding that LLMs possess extensive world knowledge and can maintain coherent personas across extended dialogues, we propose the concept of an LLM text world engine—a system where an LLM operates in dual capacity: (1) as agents following behavioral policies $\pi_{i},\pi_{j}$ derived from user personas and (2) as the environment simulating conversation dynamics and state transitions (Figure 1). This dual-capacity design benefits from recent advances in LLM-based social simulation yang2024oasis; chang2025llms; liu2024training, role-playing agent consistency liu2024roleagent; xu2025persona, multi-agent coordination shang2025united, social alignment and cognition liu2024training; liu2025synthetic; liuexploring, and controllable agent behaviors liu2025programmable, creating coherent simulation spaces for social behaviors.
@@
We formalize compatibility assessment as a reward modeling problem with an inverse RL flavor ng2000algorithms; christiano2017deep, and connect it to preference learning problem rafailov2023direct; ramesh2024group: given observed matching outcomes and fixed LLM policies approximating human behavior, we learn to extract signals from simulated interactions that predict real human preferences. This paradigm shift—from profile comparison to interaction simulation—enables us to capture the emergent properties of relationships that static features cannot represent. Theoretically, we prove that as LLM policies better approximate human behavior, compatibility predictions converge to optimal matching. Empirically, we initially validate our approach on speed dating and divorce prediction datasets.
@@
Our work makes three key contributions: (1) Paradigm and Architecture: We propose the Love First, Know Later paradigm—simulating interactions to assess compatibility—and introduce LLM text world engines that operate in dual capacity as both agents and environment. (2) Mathematical Formalization with Theoretical Guarantees: We formalize compatibility as a reward modeling problem and translate observations from relationship psychology and social science into mathematical hypotheses, bridging qualitative social insights with quantitative frameworks. Under these hypotheses, we prove convergence of our methods to optimal matching outcome. (3) Empirical Validation and Analysis: We validate feasibility through experiments on speed dating and divorce prediction, and we provide an in-depth analysis of the potential future of our paradigm. [^1] Code and data: https://anonymous.4open.science/r/test-love-first-D111
@@ 2.1 Problem Formulation
We formalize romantic compatibility prediction as a reward modeling problem christiano2017deep. Given a dataset $\mathcal{D}=\{(\pi_{i},\pi_{j},y_{ij})\}$ where $\pi_{i},\pi_{j}$ are individuals’ policies and $y_{ij}\in\{0,1\}$ indicates mutual matching, our goal is to learn a reward function $R:\Pi\times\Pi\rightarrow\mathbb{R}$ predicting human preferences (see Appendix C for MDP formulation). This builds on inverse reinforcement learning ng2000algorithms and reward modeling from human feedback christiano2017deep. One issue is we don’t have human policies $\pi_{i}^{*}$ and $\pi_{j}^{*}$. We approximate them using LLMs: $\hat{\pi}_{i}(a|s)=\text{LLM}(a|s,P_{i})$ based on persona $P_{i}$, then extract features from simulation to learn human preferences.
@@ 2.3 Learning Human Preferences via Love Observer
We propose the idea of love observer—a specialized LLM that extracts ratings from simulated interactions: Individual Participant Ratings. After interaction $\tau_{ij}$, the observer assesses each participant’s perspective independently, producing $r_{1}=\text{Observer}(\tau_{ij},p_{i})$ and $r_{2}=\text{Observer}(\tau_{ij},p_{j})$. Observer Rating. The observer also provides an external assessment $r_{3}=\text{Observer}(\tau_{ij},p_{\text{ext}})$ evaluating overall compatibility based on conversational flow, mutual engagement, and value alignment. We learn a compatibility score $R(i,j)$ by combining these ratings using a simple linear method trained on observed matching decisions. This approach is inspired by recent work on calibrating LLM judgments and learning from preference data shao2025earlier; slocum2025diverse.
@@ 3 Efficient Simulation via Critical Events
These hypotheses are grounded in relationship research showing that critical moments—conflict resolution gottman1998predicting, first-date impressions finkel2012online, value-alignment discussions ban2024neural—strongly predict outcomes, while routine interactions contribute minimally (Hypothesis 1). Moreover, trait activation theory tett2003personality and situational strength research meyer2010situational show individuals exhibit consistent behaviors in trait-relevant situations (Hypothesis 2). This motivates our two simulation modes. Speed Dating Mode tests whether LLM policies can learn human preferences for initial chemistry through brief dialogues. Critical Events Mode validates our sparse rewards hypothesis by probing fundamental compatibility through several pivotal scenarios (career conflicts, family planning)—these rare but deterministic moments reveal true compatibility. We give a theorem ensuring that improving LLM agents directly translates to better matching outcomes—justifying investment in agent refinement.
@@ 5 Discussion
(i) Personalized evolution and continuous improvement: Each user develops their own agent that improves through feedback, creating personalized compatibility predictors. Users can refine agents through preference optimization techniques—ranging from training-based methods like DPO rafailov2023direct and its variants (ramesh2024group, shao2025earlier,slocum2025diverse) to training-free approaches like In-Context DPO song-etal-2025-instantly that operate purely through prompts (one individual agent per person). As users provide more feedback and as foundation models advance liuexploring; liu2025programmable, predictions automatically improve, converging toward optimal matching.
(ii) Bidirectional interaction and transparency: Unlike black-box systems, users actively participate in the matching process through bidirectional interaction. They observe simulated interactions between their agent and potential matches, gaining transparent insights into compatibility assessments. Users provide feedback on agent behaviors and simulation quality, which refines both the agent policy and observer judgments. This continuous dialogue between user and system park2023generative; liu2024roleagent ensures persona authenticity xu2025persona while transforming matching from passive algorithm acceptance to active, collaborative exploration.
(iii) Active preference exploration: Agents probe hidden compatibility aspects through new scenarios, helping users discover what matters through simulated experiences rather than questionnaires. This draws on active learning ban2024neural and exploration principles sutton2018reinforcement. Unlike traditional methods where profiles are fixed and cannot reveal personal preferences on unasked dimensions, simulation enables dynamic exploration of unique relationship scenarios and concerns.
Despite this potential, limitations remain: (i) text-only simulations miss non-verbal cues but can be mitigated by using novel-like gesture/expression/tone description; (ii) our current dyadic modeling focuses on pairwise compatibility—extending to multi-person social dynamics (e.g., group dating, family integration, friend network compatibility) requires game-theoretic frameworks beyond simple pairwise matching. As LLMs advance in persona consistency xu2025persona; liu2024roleagent and social reasoning liu2024training, we expect this Love First, Know Later approach to unlock new possibilities wherever compatibility emerges through interaction.
@@ Appendix B Proof of Theorem 1
The policy approximation error $\epsilon$ can be reduced through online learning via preference optimization methods such as DPO rafailov2023direct and its variants. As users provide feedback on agent behaviors, $\epsilon_{t}\to 0$, ensuring convergence to optimal matching by Theorem 1.
@@ Appendix C MDP Formulation Background
We view each interaction as a multi-agent MDP $\mathcal{M}=(\mathcal{S},\mathcal{A}_{i},\mathcal{A}_{j},T,R,\gamma)$ in which two policies $\pi_{i},\pi_{j}$ act over states $s\in\mathcal{S}$ with transition dynamics $T$ and per-step rewards summarized by a compatibility signal $R$. Compatibility emerges from the returns induced by the joint policies. Since true human policies $\pi_{i}^{*},\pi_{j}^{*}$ are unobserved, we approximate them with LLM-driven policies $\hat{\pi}_{i},\hat{\pi}_{j}$ and estimate $R$ from simulated trajectories, following inverse reinforcement learning principles ng2000algorithms.
@@ D.1 Datasets
Divorce Prediction. 170 couples, 54 Likert-scale questions (1-5) covering conflict resolution, shared values, mutual understanding, and communication patterns yontem2019divorce. Outcome: married (1) vs divorced (0).

### v2

@@ 1 Introduction
Modern dating platforms primarily rely on profile similarity metrics, yet decades of relationship research show that compatibility emerges from interaction dynamics, not static attributes [finkel2012online]. We propose Love First, Know Later: instead of comparing static profiles to predict compatibility, we simulate the relationship itself first.
Recent advances show that LLMs can express consistent personality traits [park2023generative, li2024big5chat, wang2025incharacter], follow human psychological patterns [kosinski2023theory], and engage in social interactions [yang2024oasis, zhou2024sotopia, liu2024roleagent]. These models encode rich social priors from their training data, enabling them to simulate human behaviors and emotional responses. We leverage these capabilities to use LLMs as proxies for simulating romantic interactions. Building on the finding that LLMs possess extensive world knowledge and can maintain coherent personas across extended dialogues, we propose the concept of an LLM text world engine—a system where an LLM operates in dual capacity: (1) as agents following behavioral policies $\pi_{i},\pi_{j}$ derived from user personas and (2) as the environment simulating conversation dynamics and state transitions (Figure 1). This dual-capacity design benefits from recent advances in LLM-based social simulation [yang2024oasis, chang2025llms, liu2024training], role-playing agent consistency [liu2024roleagent, xu2025persona], multi-agent coordination [shang2025united], social alignment and cognition [liu2024training, liu2025synthetic, liuexploring], and controllable agent behaviors [liu2025programmable], creating coherent simulation spaces for social behaviors.
@@
We formalize compatibility assessment as a reward modeling problem with an inverse RL flavor [ng2000algorithms, christiano2017deep], and connect it to preference learning problem [rafailov2023direct, ramesh2024group]: given observed matching outcomes and fixed LLM policies approximating human behavior, we learn to extract signals from simulated interactions that predict real human preferences. This paradigm shift—from profile comparison to interaction simulation—enables us to capture the emergent properties of relationships that static features cannot represent. Theoretically, we prove that as LLM policies better approximate human behavior, compatibility predictions converge to optimal matching. Empirically, we initially validate our approach on speed dating and divorce prediction datasets.
@@
Our work makes three key contributions: (1) Paradigm and Architecture: We propose the Love First, Know Later paradigm—simulating interactions to assess compatibility—and introduce LLM text world engines that operate in dual capacity as both agents and environment. (2) Mathematical Formalization with Theoretical Guarantees: We formalize compatibility as a reward modeling problem and translate observations from relationship psychology and social science into mathematical hypotheses, bridging qualitative social insights with quantitative frameworks. Under these hypotheses, we prove convergence of our methods to optimal matching outcome. (3) Empirical Validation and Analysis: We validate feasibility through experiments on speed dating and divorce prediction, and we provide an in-depth analysis of the potential future of our paradigm.
@@ 2.1 Problem Formulation
We formalize romantic compatibility prediction as a reward modeling problem [christiano2017deep]. Given a dataset $\mathcal{D}=\{(\pi_{i},\pi_{j},y_{ij})\}$ where $\pi_{i},\pi_{j}$ are individuals’ policies and $y_{ij}\in\{0,1\}$ indicates mutual matching, our goal is to learn a reward function $R:\Pi\times\Pi\rightarrow\mathbb{R}$ predicting human preferences (see Appendix C for MDP formulation). This builds on inverse reinforcement learning [ng2000algorithms] and reward modeling from human feedback [christiano2017deep]. One issue is we don’t have human policies $\pi_{i}^{*}$ and $\pi_{j}^{*}$. We approximate them using LLMs: $\hat{\pi}_{i}(a|s)=\text{LLM}(a|s,P_{i})$ based on persona $P_{i}$, then extract features from simulation to learn human preferences.
@@ 2.3 Learning Human Preferences via Love Observer
We propose the idea of love observer—a specialized LLM that extracts ratings from simulated interactions: Individual Participant Ratings. After interaction $\tau_{ij}$, the observer assesses each participant’s perspective independently, producing $r_{1}=\text{Observer}(\tau_{ij},p_{i})$ and $r_{2}=\text{Observer}(\tau_{ij},p_{j})$. Observer Rating. The observer also provides an external assessment $r_{3}=\text{Observer}(\tau_{ij},p_{\text{ext}})$ evaluating overall compatibility based on conversational flow, mutual engagement, and value alignment. We learn a compatibility score $R(i,j)$ by combining these ratings using a simple linear method trained on observed matching decisions. This approach is inspired by recent work on calibrating LLM judgments and learning from preference data [shao2025earlier, slocum2025diverse].
@@ 3 Efficient Simulation via Critical Events
These hypotheses are grounded in relationship research showing that critical moments—conflict resolution [gottman1998predicting], first-date impressions [finkel2012online], value-alignment discussions [ban2024neural]—strongly predict outcomes, while routine interactions contribute minimally (Hypothesis 1). Moreover, trait activation theory [tett2003personality] and situational strength research [meyer2010situational] show individuals exhibit consistent behaviors in trait-relevant situations (Hypothesis 2). This motivates our two simulation modes. Speed Dating Mode tests whether LLM policies can learn human preferences for initial chemistry through brief dialogues. Critical Events Mode validates our sparse rewards hypothesis by probing fundamental compatibility through several pivotal scenarios (career conflicts, family planning)—these rare but deterministic moments reveal true compatibility. We give a theorem ensuring that improving LLM agents directly translates to better matching outcomes—justifying investment in agent refinement.
@@ 5 Discussion
(i) Personalized evolution and continuous improvement: Each user develops their own agent that improves through feedback, creating personalized compatibility predictors. Users can refine agents through preference optimization techniques—ranging from training-based methods like DPO [rafailov2023direct] and its variants ([ramesh2024group], [shao2025earlier],[slocum2025diverse]) to training-free approaches like In-Context DPO [song-etal-2025-instantly] that operate purely through prompts (one individual agent per person). As users provide more feedback and as foundation models advance [liuexploring, liu2025programmable], predictions automatically improve, converging toward optimal matching.
(ii) Bidirectional interaction and transparency: Unlike black-box systems, users actively participate in the matching process through bidirectional interaction. They observe simulated interactions between their agent and potential matches, gaining transparent insights into compatibility assessments. Users provide feedback on agent behaviors and simulation quality, which refines both the agent policy and observer judgments. This continuous dialogue between user and system [park2023generative, liu2024roleagent] ensures persona authenticity [xu2025persona] while transforming matching from passive algorithm acceptance to active, collaborative exploration.
(iii) Active preference exploration: Agents probe hidden compatibility aspects through new scenarios, helping users discover what matters through simulated experiences rather than questionnaires. This draws on active learning [ban2024neural] and exploration principles [sutton2018reinforcement]. Unlike traditional methods where profiles are fixed and cannot reveal personal preferences on unasked dimensions, simulation enables dynamic exploration of unique relationship scenarios and concerns.
Despite this potential, limitations remain: (i) text-only simulations miss non-verbal cues but can be mitigated by using novel-like gesture/expression/tone description; (ii) our current dyadic modeling focuses on pairwise compatibility—extending to multi-person social dynamics (e.g., group dating, family integration, friend network compatibility) requires game-theoretic frameworks beyond simple pairwise matching. As LLMs advance in persona consistency [xu2025persona, liu2024roleagent] and social reasoning [liu2024training], we expect this Love First, Know Later approach to unlock new possibilities wherever compatibility emerges through interaction.
@@ Appendix B Proof of Theorem 1
The policy approximation error $\epsilon$ can be reduced through online learning via preference optimization methods such as DPO [rafailov2023direct] and its variants. As users provide feedback on agent behaviors, $\epsilon_{t}\to 0$, ensuring convergence to optimal matching by Theorem 1.
@@ Appendix C MDP Formulation Background
We view each interaction as a multi-agent MDP $\mathcal{M}=(\mathcal{S},\mathcal{A}_{i},\mathcal{A}_{j},T,R,\gamma)$ in which two policies $\pi_{i},\pi_{j}$ act over states $s\in\mathcal{S}$ with transition dynamics $T$ and per-step rewards summarized by a compatibility signal $R$. Compatibility emerges from the returns induced by the joint policies. Since true human policies $\pi_{i}^{*},\pi_{j}^{*}$ are unobserved, we approximate them with LLM-driven policies $\hat{\pi}_{i},\hat{\pi}_{j}$ and estimate $R$ from simulated trajectories, following inverse reinforcement learning principles [ng2000algorithms].
@@ D.1 Datasets
Divorce Prediction. 170 couples, 54 Likert-scale questions (1-5) covering conflict resolution, shared values, mutual understanding, and communication patterns [yontem2019divorce]. Outcome: married (1) vs divorced (0).